OpenStudy (anonymous):

In 1957 the price of a McDonald’s hamburger was 43¢. In 2002, the price of a hamburger was 88¢. Assuming the price increased linearly, what was the price of a McDonald’s hamburger in 1971? a. 57¢ b. 75¢ c. 71¢ d. 44¢

OpenStudy (anonymous):

Price increase per year (in cents) = (88-43)/(2002-1957)=45/45=1 "1971 price" = "1957 price" + cents_increase_per_year * (1971-1957) = 43 + 1*14 = 57